Show complete solution. Use equation function.

1)Find the general solution to the equation (cos y + 2) dy/dx = 2x.


1
Expert's answer
2021-06-01T01:36:39-0400
"(cos y + 2) \frac{dy}{dx }= 2x\\"

By separable variable,


"(cos y + 2) dy = 2xdx\\"

Integrate through the separated equation:


"\int(cos y + 2) dy = \int 2xdx\\
siny +2y = x^2+ c"

Which is the general solution to the equation.